OpenStudy (anonymous):

Patterns with Fractions -write a rule for the sequence. then find the unknown term. 1 3/8 , 1 3/4 , 2 1/8 , __________ , 2 7/8 Rule: please help

OpenStudy (anonymous):

find 1 3/4 - 1 3/8

OpenStudy (anonymous):

how do you do that

OpenStudy (anonymous):

ok, first what is 1 3/4 as one fraction?

OpenStudy (anonymous):

18/16

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

think of the 1 as 4/4

OpenStudy (anonymous):

so it should be 4/4 + 3/4

OpenStudy (anonymous):

what is that?

OpenStudy (anonymous):

7/8

OpenStudy (anonymous):

but why 4/4 ? how did you get that

OpenStudy (anonymous):

no, you don't add the denominators

OpenStudy (anonymous):

4/4 is the same as 1

OpenStudy (anonymous):

7/4

OpenStudy (anonymous):

if the numerator equals the denominator, it equals 1

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

ok, what is 1 3/8?

OpenStudy (anonymous):

so its 3/8 + 4/4 ? right

OpenStudy (anonymous):

no, it's 3/8 + 8/8

OpenStudy (anonymous):

9/8

OpenStudy (anonymous):

no, what's 3+8

OpenStudy (anonymous):

sorry 11/8

OpenStudy (anonymous):

ok, so we have 7/4 - 11/8

OpenStudy (anonymous):

we need to find a common denominator

OpenStudy (anonymous):

what times 4 equals 8?

OpenStudy (anonymous):

its 2

OpenStudy (anonymous):

so we multiply 7/4 by 2

OpenStudy (anonymous):

this gives us 14/8

OpenStudy (anonymous):

what's (14-11)/8?

OpenStudy (anonymous):

14-11=?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

3/8

OpenStudy (anonymous):

That's the difference between each term

OpenStudy (anonymous):

we know that the rule is that you add 3/8 to get the next term

OpenStudy (anonymous):

17/8 + 3/8 =

OpenStudy (anonymous):

solve that to get the missing term

OpenStudy (anonymous):

20/8

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

now you have to reduce it

OpenStudy (anonymous):

2 1/5 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

The 2 is right though

OpenStudy (anonymous):

20-16=?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

ya so we get 2 4/8

OpenStudy (anonymous):

The 4/8 can be reduced again to 1/2

OpenStudy (anonymous):

so the answer is 2 1/2

OpenStudy (anonymous):

oh okay Thank you Very much :)

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

wait how did you get " 17/8 + 3/8 ="

OpenStudy (anonymous):

2 1/8 is 17/8 and 3/8 is what we needed to add to that term to get the unknown

OpenStudy (anonymous):